OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 25? x squared over 25 ±square root of the quantity x plus 25 ±5 times the square root of x 1 over quantity x squared minus 25

OpenStudy (anonymous):

if i did the math correctly the answer is D

OpenStudy (anonymous):

optioN B

OpenStudy (anonymous):

what did u get @Morgan27, i was drawn between B and D

OpenStudy (anonymous):

I got b

OpenStudy (anonymous):

oh ok, haha sorry, @zaibali.qasmi is correct

OpenStudy (anonymous):

Lol its okay :)

OpenStudy (anonymous):

I wasnt sure about it either

OpenStudy (anonymous):

yea, those problems are tricky

OpenStudy (anonymous):

Do you think you could help me with one more? :)

OpenStudy (anonymous):

sure!

OpenStudy (anonymous):

given the function f(x)=6(x+2)-3 solve for the inverse function when x=21 A.-45 B.0 C.2 D.63

OpenStudy (anonymous):

ok Step 1: Replace f(x) with y. Step 2: Interchange x with y. Step 3: Solve for y. If this y is a function, it is the inverse of the original function. If this y is NOT a function, there is no inverse. Step 4: If f has an inverse function, then replace y with f inverse. ok lets walk thru this problem first lets rearrange the problem. Given the function f(x) = 6(x+2) − 3 f(x)= 6x+12 -3 = 6x+9 Step 1 y=6x+9 step 2 x=6y+9 step3 can you solve for y?

OpenStudy (anonymous):

y=1/6x-3/2 ?

OpenStudy (anonymous):

yes!